OpenStudy (anonymous):

Solve the following equation for x: ln(2-x)+ ln(3x)=4

myininaya (myininaya):

ln(x)+ln(y)=ln(xy)

myininaya (myininaya):

ln[(2-x)(3x)]=4

myininaya (myininaya):

to get rid of the natural log base e both sides

myininaya (myininaya):

e^(ln[(2-x)(3x)]=e^4

myininaya (myininaya):

so we have (2-x)(3x)=e^4

myininaya (myininaya):

6x-3x^2-e^4=0

myininaya (myininaya):

we can use the quadratic formula now!

myininaya (myininaya):

since we have -3x^2+6x-e^4=0, we have a=-3, b=6, c=-e^4

myininaya (myininaya):

so x=(-6 @ sqrt(36-4*(-3)*(-e^4)))/(-6) where @ symbolizes + or -

myininaya (myininaya):

make sure your x is positve because you cannot have natural log of a negative number